# Structural-Spectral Graph Convolution with Evidential Edge Learning for Hyperspectral Image Clustering

### v1

@@ Abstract
Hyperspectral image (HSI) clustering assigns similar pixels to the same class without any annotations, which is an important yet challenging task. For large-scale HSIs, most methods rely on superpixel segmentation and perform superpixel-level clustering based on graph neural networks (GNNs). However, existing GNNs cannot fully exploit the spectral information of the input HSI, and the inaccurate superpixel topological graph may lead to the confusion of different class semantics during information aggregation. To address these challenges, we first propose a structural-spectral graph convolutional operator (SSGCO) tailored for graph-structured HSI superpixels to improve their representation quality through the co-extraction of spatial and spectral features. Second, we propose an evidence-guided adaptive edge learning (EGAEL) module that adaptively predicts and refines edge weights in the superpixel topological graph. We integrate the proposed method into a contrastive learning framework to achieve clustering, where representation learning and clustering are simultaneously conducted. Experiments demonstrate that the proposed method improves clustering accuracy by 2.61%, 6.06%, 4.96% and 3.15% over the best compared methods on four HSI datasets. Our code is available at https://github.com/jhqi/SSGCO-EGAEL.
Index Terms:
Hyperspectral image, clustering, superpixel, graph neural networks, contrastive learning.
@@ I Introduction
Hyperspectral image (HSI) captures the reflectance of land covers across hundreds of contiguous narrow bands, providing rich and precise spectral information. Classifying the pixels of HSI into different categories is a fundamental task [1, 2, 3, 4, 5, 6], as it describes the spatial distribution of different land covers. This technique is widely applied in various fields, such as precision agriculture [7], environmental monitoring [8], mineral exploration [9], and urban planning [10, 11, 12].
@@
Due to the high cost of obtaining HSI pixel annotations, unsupervised clustering methods that explore the relationships between pixels without labels have gained increasing attention. Traditional clustering methods, e.g., K-means [13], fuzzy c-means [14] and subspace clustering [15, 16], have been applied to HSI clustering, but with relatively poor clustering performance. This is because those methods usually perform clustering based on the raw feature of HSI, which is susceptible to spectral noise of HSI caused by the illumination condition or the viewing angle [17]. To this end, research has increasingly turned to deep clustering methods [18, 19, 20], which leverage deep neural networks to learn discriminative feature representations, achieving effective clustering results under end-to-end training. However, the massive number of pixels in HSI may lead to high computational cost. Consequently, many studies employed superpixel segmentation techniques [21, 22, 23] and transformed the clustering task from the pixel level to the superpixel level to reduce the computational cost. Superpixel-level clustering has currently become the mainstream approach for large-scale HSI clustering [24, 25, 26, 27, 28].
Given the irregular shapes and clear spatial distribution, superpixels are naturally modeled as graph-structured data, where each superpixel acts as a node and the topological graph is constructed based on their spatial adjacency or spectral similarity. With the powerful capability of information aggregation, graph neural networks (GNNs) have excelled in graph learning tasks [29, 30, 31]. Consequently, several studies have employed GNNs for superpixel representation learning to incorporate topological information into superpixel embeddings and then conducted GNN-based clustering, which can be mainly categorized into two types. The first type is based on graph autoencoders (GAEs) [32], which typically reconstruct the graph adjacency matrix from node embeddings, and introduce an auxiliary distribution loss [18] for self-supervised training [33, 34, 35, 36, 37]. The second type utilizes contrastive clustering frameworks [38, 39, 40], conducting alignment or contrast between samples or prototypes to improve the discriminability of superpixel representations [19, 20, 41]. Although the above methods have achieved promising results, two critical limitations still exist.
@@
The rest of this paper is organized as follows. Section II introduces the adopted data preprocessing approaches, including superpixel segmentation and graph construction in HSIs. Section III presents the proposed method in detail. In Section IV, we provide experimental results and analyses on four HSI datasets. Finally, we conclude this work in Section V.
Figure 2: The diagram of the proposed method: (a) the contrastive clustering framework consists of an online network $f(\cdot)$ and a target network $f^{\prime}(\cdot)$, which take superpixels and augmented views as inputs. Neighborhood alignment (NA) and prototype contrast (PC) are adopted as training objectives; (b) Structural-spectral graph convolutional operator (SSGCO) simultaneously performs 1D convolution (1D-Conv) and graph convolution in each layer to extract both spatial and spectral features of superpixels; (c) In evidence guided adaptive edge learning (EGAEL), each edge is represented by concatenating the soft assignments of its two endpoint nodes, and input into an MLP to get the predicted edge weight $w^{pre}$. The empirical edge weight $w^{emp}$ computed with node clustering confidence and similarity is adopted as guidance. The adjacency matrix is momentum updated based on $w^{pre}$.
@@ II Data Preprocessing
Then, we apply PCA on raw pixel features again, and retain only $d$ components to further reduce the data dimensionality. Let $p_{i}$ denotes the $i$-th pixel, $sp_{j}$ denotes the $j$-th superpixel, $\bm{x}^{p}_{i}\in\mathbb{R}^{d}$ and $\bm{x}^{sp}_{j}\in\mathbb{R}^{d}$ are their feature vectors respectively, following:
$\bm{x}^{sp}_{j}=\frac{1}{|sp_{j}|}\sum_{p_{i}\in sp_{j}}{\bm{x}^{p}_{i}},$ (1)
where $|sp_{j}|$ denotes the number of pixels in $sp_{j}$. For the sake of simplicity, we abbreviate $\bm{x}^{sp}_{j}$ as $\bm{x}_{j}$. Thus, features vectors of all the superpixels are represented as $\mathbf{X}=[\bm{x}_{1};...;\bm{x}_{M}]\in\mathbb{R}^{M\times d}$, where each row corresponds to a superpixel.
@@
where $\mathbf{I}_{M}\in\mathbb{R}^{M\times M}$ is the identity matrix, $\widetilde{\mathbf{D}}$ is the degree matrix of the adjacency matrix with self-loop, and $\widehat{\mathbf{A}}$ is the symmetrically normalized adjacency matrix.
@@ III-A Contrastive Clustering Framework
Given one prototype, the remaining $K-1$ prototypes are definitely negative examples. We construct a contrastive loss [43] to encourage the alignment between different views of the same prototype, and increase the inter-cluster distance, i.e.,
@@ III-B Structural-Spectral Graph Convolutional Operator
While existing GNNs excel at leveraging spatial information of HSI based on the superpixel topological graph, they often inadequately extract deep spectral features. Their reliance on projective transformations (i.e., linear layers) limits the capability to deal with the inherent continuity and redundancy of HSI spectra. To address this limitation, we propose a novel structural-spectral graph convolutional operator (SSGCO) to enhance the superpixel representation learning. SSGCO enriches the information aggregation capability of GNNs by infusing dedicated 1D convolutional kernels that target high-order spectral feature extraction, thus simultaneously learning discriminative spectral features and aggregating spatial neighborhood information at the superpixel level.
Based on the standard multi-layer GCN, we additionally introduce 1D convolution along the spectral dimension at each layer to enhance its ability to extract spectral features, as illustrated in Fig. 2b. Then, SSGCO is formally defined as:
$\mathbf{H}_{(l)}=\phi(\widehat{\mathbf{A}}(\mathbf{H}_{(l-1)}\circledast_{\scriptscriptstyle\mathrm{1D}}\mathbf{W}^{c}_{(l)})\mathbf{W}^{g}_{(l)}),$ (12)
@@ III-C Evidence-guided Adaptive Edge Learning
where ||||| | denotes the concatenating operation. Then, we employ a two-layer MLP $h(\cdot)$ to predict the new edge weight $w_{u,v}^{pre}$:
@@ III-D Training Procedure
TABLE II: Overview of Indian Pines, Pavia University, Botswana and Trento.
Dataset     Sensor Resolution #Spectral Band #Class #Labeled Pixel
Indian Pines     AVIRIS 145 $\times$ 145 200 16 10249
Pavia University     ROSIS 610 $\times$ 340 103 9 42776
Botswana     Hyperion 1476 $\times$ 256 145 14 3248
Trento     AISA Eagle 600 $\times$ 166 63 6 30214
@@ IV-A Datasets
We evaluated our method on four public HSI datasets, including Indian Pines, Pavia University, Botswana and Trento. An overview of all the datasets is shown in Table II.
@@
TABLE III: Clustering Performance on HSI Datase, Indian Pines, Pavia University, Bowana and Trento are Abbreviated to IP, PU, BO and TR, $\bullet$/$\circ$ Indicates whether Our Method is Statistically Superior/Inferior to the Compared Method According to the Pairwise $t$-Test at 0.05 Significance Level.
@@ IV-B Experiment Setup
1) Implementation Details: The input HSI is standardized across each spectral band. To decrease computational costs, PCA is adopted to reduce the number of bands to 40, 20, 25 and 40 on Indian Pines, Pavia University, Botswana and Trento, respectively. In the contrastive clustering framework, the hidden size of the predictor $g(\cdot)$ is set to 512. The momentum coefficient $m$ for updating $f^{\prime}(\cdot)$ is set to 0.99. For neighborhood alignment, $\sigma$ is set to $1\times 10^{-3}$. For prototype contrast, $\tau$ is set to 0.7. In the structural-spectral graph convolutional operator, the kernel size of 1D convolution in the first layer is set to 7, with subsequent layers decreasing by 2, but not smaller than 3. The output channels in the first layer is set to 16, with subsequent layers multiplying by 2, but not larger than 64. In the evidence-guided adaptive edge learning, $h(\cdot)$ consists of a two-layer MLP with the hidden size of $K$. We used the SGD optimizer with the learning rate, weight decay, and momentum coefficient set to 0.05, $5\times 10^{-4}$, and 0.9, respectively. In addition, we adopted cosine annealing and set the learning rate of $g(\cdot)$ 10×10$\times$10 × as that of other modules following [39]. We repeated all experiments five times and reported the mean and standard deviation of the results. We carried out all experiments in a Linux server with an AMD EPYC 7352 CPU and an NVIDIA RTX 4090 GPU.
@@
K-means [13] is a traditional centroid-based clustering algorithm with K-means++ initialization.
@@
Figure 4: Clustering maps on Pavia University: (a) ground truth, (b) K-means, (c) SSC, (d) NCSC, (e) SGLSC, (f) DGAE, (g) S3-ULDA, (h) S2DL, (i) SDST, (j) SPGCC, (k) SAPC, (l) Ours.
@@ IV-C Comparison with State-of-the-Art Methods
We report the clustering performance of our method and other compared methods on Indian Pines, Pavia University, Botswana, and Trento in Table III. For each metric, the best value is marked in bold, while the second-best value is underlined.
According to the quantitative clustering metrics, it could be observed that our method outperforms other compared methods with a large margin. For instance, our method improves the ACC by 2.61%, 6.06%, 4.96% and 3.15% on the four datasets. Notably, on Pavia University, ACC, $\mathcal{K}$, ARI and F1 obtained by our method outperform the second-best value by 6.06%, 7.47%, 8.22% and 1.13%, respectively. Overall, on Botswana and Trento, our method achieves the best scores on all metrics; on Indian Pines, it achieves the best scores on five metrics and the second-best on three metrics; on Pavia University, it achieves the best scores on seven metrics and the second-best on one metric. Based on the results of pairwise $t$-Test at 0.05 significance level, our method significantly outperforms the compared methods in 91.3% (292/320) of cases.
@@
Moreover, clustering map visualizations of different methods on Indian Pines and Pavia University are reported in Figs. 3 and 4. According to the visualizations, it is evident that our method demonstrates superior clustering for large-scale and continuously distributed land covers, e.g., dark green and dark yellow areas in Indian Pines, light green and light yellow areas in Pavia University. This advantage primarily stems from our method performing superpixel-level clustering and adopting the edge learning to construct a more accurate adjacency matrix for superpixels.
TABLE IV: Running Time of Different Methods (s).
Method     Indian Pines Pavia University Botswana Trento     Avg.
K-means     22.6 113.7 273.9 23.7     108.5
SSC     20.3 16.4 22.3 9.7     17.2
NCSC     77.5 6574.4 8559.3 2900.6     4528.0
SGLSC     108.0 195.6 368.5 173.0     211.3
DGAE     25.5 202.0 12.9 44.4     71.2
S3-ULDA     11.5 374.5 2792.4 88.9     816.8
S2DL     12.5 544.9 2960.6 65.7     895.9
SDST     296.9 229.7 169.0 129.2     206.2
SPGCC     85.7 619.9 1129.2 437.3     568.0
SAPC     11.1 21.8 51.6 13.2     24.4
Ours     35.0 58.5 51.3 113.0     64.5
TABLE V: Optimal Hyperparameters.
@@
IV-D Running Time Analysis
IV-E Hyperparameters Analysis
The proposed method mainly requires five hyperparameters: number of superpixels ($M$), number of network layers ($L$), weight of $\mathcal{L}_{PC}$ ($\alpha$), weight of $\mathcal{L}_{E}$ ($\beta$) and momentum coefficient of edge weight updating ($\gamma$). We have tuned these hyperparameters for different datasets, and the adopted values are reported in Table V.
@@ IV-F Further Analysis of Structural-Spectral Graph Convolutional Operator
According to the experimental results presented in Table VI, the proposed SSGCO outperforms all other variant networks with a large margin. In addition, we have identified two observations.
Figure 5: The influence of different numbers of superpixels $M$ and different numbers of network layers $L$ on: (a) Indian Pines, (b) Pavia University, (c) Botswana and (d) Trento.
IV-G Further Analysis of Evidence Guided Adaptive Edge Learning
TABLE VI: Clustering Performance of Variant Networks.

### v2

@@ Abstract
Hyperspectral image (HSI) clustering groups pixels into clusters without labeled data, which is an important yet challenging task. For large-scale HSIs, most methods rely on superpixel segmentation and perform superpixel-level clustering based on graph neural networks (GNNs). However, existing GNNs cannot fully exploit the spectral information of the input HSI, and the inaccurate superpixel topological graph may lead to the confusion of different class semantics during information aggregation. To address these challenges, we first propose a structural-spectral graph convolutional operator (SSGCO) tailored for graph-structured HSI superpixels to improve their representation quality through the co-extraction of spatial and spectral features. Second, we propose an evidence-guided adaptive edge learning (EGAEL) module that adaptively predicts and refines edge weights in the superpixel topological graph. We integrate the proposed method into a contrastive learning framework to achieve clustering, where representation learning and clustering are simultaneously conducted. Experiments demonstrate that the proposed method improves clustering accuracy by 2.61%, 6.06%, 4.96% and 3.15% over the best compared methods on four HSI datasets. Our code is available at https://github.com/jhqi/SSGCO-EGAEL.
@@ I Introduction
Hyperspectral image (HSI) captures the reflectance of land covers across hundreds of contiguous narrow bands, providing rich and precise spectral information. Classifying the pixels of HSI into different categories is a fundamental task [1, 2, 3, 4, 5, 6, 7], as it describes the spatial distribution of different land covers. This technique is widely applied in various fields, such as precision agriculture [8], environmental monitoring [9], mineral exploration [10], and urban planning [11].
@@
Due to the high cost of obtaining HSI pixel annotations, unsupervised clustering methods that explore the relationships between pixels without labels have gained increasing attention. Traditional clustering methods, e.g., K-means [12], fuzzy c-means [13] and subspace clustering [14], have been applied to HSI clustering, but with relatively poor clustering performance. This is because those methods usually perform clustering based on the raw feature of HSI, which is susceptible to spectral noise of HSI caused by the illumination condition or the viewing angle [15]. To this end, research has increasingly turned to deep clustering methods [16, 17, 18], which leverage deep neural networks to learn discriminative feature representations, achieving effective clustering results under end-to-end training. However, the massive number of pixels in HSI may lead to high computational cost. Consequently, many studies employed superpixel segmentation techniques [19, 20, 21] and transformed the clustering task from the pixel level to the superpixel level to reduce the computational cost. Superpixel-level clustering has currently become the mainstream approach for large-scale HSI clustering [22, 23, 24, 25, 26].
Given the irregular shapes and clear spatial distribution, superpixels are naturally modeled as graph-structured data, where each superpixel acts as a node and the topological graph is constructed based on their spatial adjacency or spectral similarity. With the powerful capability of information aggregation, graph neural networks (GNNs) have excelled in graph learning tasks [27, 28]. Consequently, several studies have employed GNNs for superpixel representation learning to incorporate topological information into superpixel embeddings and then conducted GNN-based clustering, which can be mainly categorized into two types. The first type is based on graph autoencoders (GAEs) [29], which typically reconstruct the graph adjacency matrix from node embeddings, and introduce an auxiliary distribution loss [16] for self-supervised training [30, 31, 32, 33, 34]. The second type utilizes contrastive clustering frameworks [35, 36, 37], conducting alignment or contrast between samples or prototypes to improve the discriminability of superpixel representations [17, 18, 38]. Although the above methods have achieved promising results, two critical limitations still exist.
@@
Figure 2: The diagram of the proposed method: (a) the contrastive clustering framework consists of an online network $f(\cdot)$ and a target network $f^{\prime}(\cdot)$, which take superpixels and augmented views as inputs. Neighborhood alignment (NA) and prototype contrast (PC) are adopted as training objectives; (b) Structural-spectral graph convolutional operator (SSGCO) sequentially performs 1D convolution (1D-Conv) and graph convolution in each layer to extract both spatial and spectral features of superpixels; (c) In evidence guided adaptive edge learning (EGAEL), each edge is represented by concatenating the soft assignments of its two endpoint nodes, and input into an MLP to get the predicted edge weight $w^{pre}$. The empirical edge weight $w^{emp}$ computed with node clustering confidence and similarity is adopted as guidance. The adjacency matrix is momentum updated based on $w^{pre}$.
@@ II Data Preprocessing
Then, we apply PCA on raw pixel features again, and retain only $d$ components to further reduce the data dimensionality [33]. For the $i$-th pixel $p_{i}$ and the $j$-th superpixel $sp_{j}$, their feature vectors are denoted as $\bm{x}^{p}_{i}\in\mathbb{R}^{d}$ and $\bm{x}_{j}\in\mathbb{R}^{d}$, following:
$\bm{x}_{j}=\frac{1}{|sp_{j}|}\sum_{p_{i}\in sp_{j}}{\bm{x}^{p}_{i}},$ (1)
where $|sp_{j}|$ denotes the number of pixels in $sp_{j}$. Thus, features vectors of all the superpixels are represented as $\mathbf{X}=[\bm{x}_{1};...;\bm{x}_{M}]\in\mathbb{R}^{M\times d}$, where each row corresponds to a superpixel.
@@
where $\mathbf{I}_{M}\in\mathbb{R}^{M\times M}$ is the identity matrix, $\widetilde{\mathbf{D}}$ is the degree matrix with self-loop, and $\widehat{\mathbf{A}}$ is the symmetrically normalized adjacency matrix.
@@ III-A Contrastive Clustering Framework
Given one prototype, the remaining K1K-1 prototypes are definitely negative examples. We construct a contrastive loss [40] to encourage the alignment between different views of the same prototype, and increase the inter-cluster distance, i.e.,
@@ III-B Structural-Spectral Graph Convolutional Operator
While existing GNNs excel at leveraging spatial information of HSI based on the superpixel topological graph, they often inadequately extract deep spectral features. Their reliance on projective transformations (i.e., linear layers) limits the capability to deal with the inherent continuity and redundancy of HSI spectra. To address this limitation, we propose a novel structural-spectral graph convolutional operator (SSGCO) to enhance the superpixel representation learning. SSGCO enriches the information aggregation capability of GNNs by infusing dedicated 1D convolutional kernels that target high-order spectral feature extraction.
Based on the standard multi-layer GCN, each SSGCO layer sequentially applies 1D convolution to extract deep spectral patterns, followed by graph convolution to aggregate spatial information, as illustrated in Fig. 2b. SSGCO is formally defined as:
[…]
@@ III-C Evidence-guided Adaptive Edge Learning
where $||$ denotes the concatenating operation. Then, we employ a two-layer MLP $h(\cdot)$ to predict the new edge weight $w_{u,v}^{pre}$:
@@ IV-A Datasets
We evaluated our method on four public HSI datasets, including Indian Pines, Pavia University, Botswana and Trento.
@@
TABLE II: Clustering Performance on HSI Datasets, Indian Pines, Pavia University, Botswana and Trento are Abbreviated to IP, PU, BO and TR, $\bullet$$\circ$ Indicates whether Our Method is Statistically Superior/Inferior to the Compared Method According to the Pairwise $t$-Test at 0.05 Significance Level.
@@ IV-B Experiment Setup
1) Implementation Details: The input HSI is standardized across each spectral band. To decrease computational costs, PCA is adopted to reduce the number of bands to 40, 20, 25 and 40 on Indian Pines, Pavia University, Botswana and Trento, respectively. In the contrastive clustering framework, the hidden size of the predictor $g(\cdot)$ is set to 512. The momentum coefficient $m$ for updating $f^{\prime}(\cdot)$ is set to 0.99. For neighborhood alignment, $\sigma$ is set to $1\times 10^{-3}$. For prototype contrast, $\tau$ is set to 0.7. In the structural-spectral graph convolutional operator, the kernel size of 1D convolution in the first layer is set to 7, with subsequent layers decreasing by 2, but not smaller than 3. The output channels in the first layer is set to 8, with subsequent layers multiplying by 2, but not larger than 64. FC layers in graph convolution keep output dimensions equal to input dimensions. In the evidence-guided adaptive edge learning, $h(\cdot)$ consists of a two-layer MLP with the hidden size of $K$. We used the SGD optimizer with the learning rate, weight decay, and momentum coefficient set to 0.05, $5\times 10^{-4}$, and 0.9, respectively. In addition, we adopted cosine annealing and set the learning rate of $g(\cdot)$ $10\times$ as that of other modules following [36]. We repeated all experiments five times and reported the mean and standard deviation of the results. We carried out all experiments in a Linux server with an AMD EPYC 7352 CPU and an NVIDIA RTX 4090 GPU.
@@
K-means [12] is a traditional centroid-based clustering algorithm with “K-means++” initialization.
@@
Figure 4: The influence of different numbers of superpixels ($M$) and different numbers of network layers ($L$) on: (a) Indian Pines, (b) Pavia University, (c) Botswana and (d) Trento.
@@ IV-C Comparison with State-of-the-Art Methods
We report the clustering performance of our method (SSGCO-EGAEL) and other compared methods on Indian Pines, Pavia University, Botswana, and Trento in Table II. For each metric, the best value is marked in bold, while the second-best value is underlined.
According to the quantitative clustering metrics, it could be observed that our method outperforms other compared methods with a large margin. For instance, our method improves the ACC by 2.61%, 6.06%, 4.96% and 3.15% on the four datasets. Notably, on Pavia University, ACC, $\mathcal{K}$, ARI and F1 obtained by our method outperform the second-best value by 6.06%, 7.47%, 8.22% and 1.13%, respectively. Overall, on Botswana and Trento, our method achieves the best scores on all metrics; on Indian Pines, it achieves the best scores on five metrics and the second-best on three metrics; on Pavia University, it achieves the best scores on seven metrics and the second-best on one metric. Based on the results of pairwise $t$-Test at 0.05 significance level, our method significantly outperforms the compared methods in 90.9% (291/320) cases.
@@
Moreover, the clustering map visualization of different methods on Indian Pines is reported in Fig. 3. (Clustering map visualizations of other datasets are reported in the supplementary material.) According to the visualizations, it is evident that our method demonstrates superior clustering for large-scale and continuously distributed land covers, e.g., green and yellow areas in Indian Pines. This advantage primarily stems from our method performing superpixel-level clustering and adopting the edge learning to construct a more accurate adjacency matrix for superpixels.
TABLE III: Optimal Hyperparameters.
@@
TABLE IV: Running Time of Different Methods (s), Indian Pines, Pavia University, Botswana and Trento are Abbreviated to IP, PU, BO and TR.
Method     IP PU BO TR     Avg.     Rk.
K-means     22.6 113.7 273.9 23.7     108.5     5
SSC     20.3 16.4 22.3 9.7     17.2     1
NCSC     77.5 6574.4 8559.3 2900.6     4528.0     11
SGLSC     108.0 195.6 368.5 173.0     211.3     7
DGAE     25.5 202.0 12.9 44.4     71.2     4
S3-ULDA     11.5 374.5 2792.4 88.9     816.8     9
S2DL     12.5 544.9 2960.6 65.7     895.9     10
SDST     296.9 229.7 169.0 129.2     206.2     6
SPGCC     85.7 619.9 1129.2 437.3     568.0     8
SAPC     11.1 21.8 51.6 13.2     24.4     2
Ours     35.0 58.5 51.3 113.0     64.5     3
IV-D Hyperparameters Analysis
The proposed method mainly requires five hyperparameters: number of superpixels ($M$), number of network layers ($L$), weight of $\mathcal{L}_{PC}$ ($\alpha$), weight of $\mathcal{L}_{E}$ ($\beta$) and momentum coefficient of edge weight updating ($\gamma$). We have tuned these hyperparameters for different datasets, and the adopted values are reported in Table III.
@@ IV-D Hyperparameters Analysis
IV-E Running Time Analysis
TABLE V: Clustering Performance of Variant Networks.